# A predictive modular approach to constraint satisfaction under uncertainty - with application to glycosylation in continuous monoclonal antibody biosimilar production

### v1

@@ 6 Conclusion and Discussion
In this paper we proposed a universal modular approach to constraint handling, based on a predictive filter combined with constraint margin adaptation and constraint violation cost monitor to minimize the cost of soft constraint violations. The main advantage of the proposed method is that, as a universal module, it can be easily applied to a system with arbitrary controllers and can ensure constraint satisfaction with minimal modification of the desired closed-loop performance. This provides a simple approach to enable systems to complete tasks using arbitrary controllers or optimizers without significant economic cost induced by high-cost soft constraint violations, which is of increasing importance due to the increased interest in applying learning-based algorithms in bioproduction systems. By constantly monitoring the constraint violation cost, the proposed method can adjust optimal constraint margins according to the changes in the size and characteristics of uncertainties. Both the learning of the optimal constraint margin and the detection of the changes in uncertainties are directly from measurement data, no prior knowledge of the uncertainties is required. We illustrated the potential of the proposed method using a realistic simulation case study of glycosylation constraint satisfaction in continuous mAb biosimilar production based on a realistic metabolic network and glycosylation models under uncertainty, including unmeasured disturbances, involving 23 extracellular metabolites and 126 reactions. We also demonstrated the advantages of feedback control of glycosylation under uncertainty. The efficiency of the constraint margin adaptation and constraint satisfaction, while largely maintaining the desired closed-loop behaviour, demonstrates the potential of the proposed method for application in complex control and optimization systems that do not explicitly handle constraints.

### v3

@@ 6 Conclusion and Discussion
In this paper we proposed a universal modular approach to constraint handling, based on a predictive filter combined with constraint margin adaptation and constraint violation cost monitoring to minimize the cost of soft constraint violations. The main advantage of the proposed method is that, as a universal module, it can be easily applied to a system with arbitrary controllers and can ensure constraint satisfaction with minimal modification of the desired closed-loop performance. This provides a simple approach to enable systems to complete tasks using arbitrary controllers or optimizers without significant economic cost induced by high-cost soft constraint violations. By constantly monitoring the constraint violation cost, the proposed method can adjust optimal constraint margins according to changes in the size and characteristics of uncertainties. Both the learning of the optimal constraint margin and the detection of the changes in uncertainties are directly from real-time data, and no prior knowledge of the uncertainties is required. We illustrated the potential of the proposed method using a realistic simulation case study of glycosylation constraint satisfaction in continuous mAb biosimilar production based on a realistic metabolic network and glycosylation models involving 23 extracellular metabolites and 126 reactions. The method resulted in a more than 60% reduction in the cost due to constraint violations as compared to the case without the proposed constraint handling method. The efficiency of the constraint margin adaptation and constraint satisfaction, while largely maintaining the desired closed-loop behaviour, demonstrates the potential of the proposed method for application in complex control and optimization systems that do not explicitly handle constraints.